Calculate ΔG0 at 298 K for the reaction below.

Fe2O3(s) + 13CO(g) > 2Fe(CO)5(g) + 3CO2(g)


ΔH0f (kJ/mol) -824.2 -110.5 -733.8 -393.5

S0 (J/mol•K) 87.4 197.6 445.2 213.6


1
Expert's answer
2022-05-31T09:21:03-0400

∆H of the reaction = 3*∆H(CO2) + 2*∆H(Fe(CO)5) - 13*∆H(CO) - ∆H(Fe2O3):


∆H of the reaction = 3 *(-393,5)+ 2 *(-733,8) - (13 *(-110,5)) - (-824,2) = -387,4 kJ


∆S of the reaction = 3*∆S(CO2) + 2*∆S(Fe(CO)5) - 13*∆S(CO) - ∆S(Fe2O3):


∆S of the reaction = 3 * 213,6 + 2 * 445,2 - (13 * 197,6) - (87,4) = -1125 J = - 1,125 kJ


∆G = ∆H - T * ∆S


∆G = -387,4 - 298 * (-1,125) = -52,15 kJ


ANSWER: -52,15 kJ